Question 2 Answer the following questions for the terms provided. Does the term get larger or smaller if z increases? 広
Answer #1

1) Increase in x increases the value of term because x increase the numerator.

If value of numerator increases keeping denominator constant then value of term increases and vice-versa.

Example; let x = 1 then value of term = (4 + 1)/5 = 5/5 = 1

If value of x increases to 6 then value of term = (4 + 6)/5 = 10/5 = 2

2) Increase in x decreases the value of term because denominator increases with increase in x.

If value of denominator increases keeping numerator constant then value of term decreases and vice-versa.

Example; let x = 0 then value of term = 1/(10 + 0) = 1/10 = 0.1

If value of x increases to 10 then value of term = 1/(10 + 10) = 1/20 = 0.05

3) Increase in x increases the value of term as increase in x decreases denominator.

Example; let x = 0 then value of term = 1/(10 - 0) = 1/10 = 0.1

If value of x increases to 5 then value of term = 1/(10 - 5) = 1/5 = 0.2

4) Increase in x decreases the value of term.

Example; let x = 1 then value of term = 1/10(1) = 1/10 = 0.1

If value of x increases to 2 then value of term = 1/10(2) = 1/20 = 0.05